### v3

# Conformal Prediction with Upper and Lower Bound Models

Abstract
Decision makers routinely use optimization technology to plan and operate complex systems like global supply chains or power grids. In that context, practitioners face the challenge of assessing the optimality gap of a given solution, i.e., how optimal or sub-optimal said solution is. While constrained optimization algorithms generally provide such information in the form of so-called dual bounds, these bounds are often too loose to be informative. Therefore, decision makers must evaluate solution quality using their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions. This paper addresses that challenge by providing estimates of solution quality in the form of prediction intervals, obtained via conformal prediction techniques. To that end, the paper introduces a general framework for conformal prediction with upper and lower bound predictors, which explicitly addresses the potential heteroskedasticity of either predictor via a model selection step. The paper also proposes a principled mechanism to avoid paradoxical miscoverage in cases where upper/lower predictors are both highly accurate. Finally, numerical experiments on large-scale problems found in industry demonstrate that the proposed approach achieves substantial reduction in interval width over several baseline methods.
1 Introduction
Optimization technology is instrumental in operating complex systems efficiently, from supply chains, logistics and manufacturing to healthcare and power systems. Nevertheless, it is often intractable to obtain optimal solutions to large-scale, nonlinear and non-convex problems found in industry. This limitation has motivated the use of approximate and heuristic algorithms, which are fast but may return sub-optimal solutions. As a consequence, decision makers face the hurdle of assessing whether a given solution is optimal enough, or whether substantial improvements are possible. This assessment requires computing dual bounds, i.e., mathematical certificates of solution quality, which are typically obtained by solving relaxations Boyd and Vandenberghe (2004); Geoffrion (2009) or by branch-and-bound algorithms Wolsey and Nemhauser (1999). However, in real-life applications, these dual bounds may be too loose to be informative, thus leaving practitioners to evaluate solution quality by relying on their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions.
To address this challenge, this paper proposes a data-driven approach that evaluates the sub-optimality of a solution by leveraging uncertainty quantification (UQ) and conformal prediction (CP) techniques. Namely, the paper proposes to provide probabilistic guarantees for solution quality in the form of prediction intervals. To the best of the authors’ knowledge, this paper is the first to propose the use of CP to estimate solution sub-optimality in an optimization setting. In doing so, the paper introduces CPUL, a general framework for CP with upper and lower bounds. This contrasts with existing approaches from the optimization literature, which almost exclusively rely on computing valid, but possibly loose, dual bounds. The paper’s setting also contrasts with traditional CP settings that rely on approximate predictors, e.g., mean predictors Vovk et al. (2005) or quantile regression models Romano et al. (2019) which, in the present setting, may not yield valid lower/upper bounds.
The paper’s contributions are summarized as follows: (1) it presents a data-driven methodology to evaluate the optimality of solutions to optimization problems using uncertainty quantification techniques, (2) it introduces CPUL, a general framework for CP with upper and lower bounds, with a focus on capturing the heteroskedasticity of lower and upper bound residuals, (3) it addresses the paradox of miscoverage depsite high-quality lower and upper bounds by proposing an additional mechanism, Optimal Minimum Length Threshold (OMLT), (4) it conducts numerical experiments on large-scale economic dispatch problems used in real-time electricity markets.
1.1 Problem statement
Motivating Example
Consider a parametric optimization problem of the form
$\Phi(x)=\min_{s}\quad$ $f_{x}(s)$ (1a)
s.t. $g_{x}(s)\leq 0,$ (1b)
where $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ denotes the instance’s parameters, $s\in\mathbb{R}^{n}$ are decision variables, $f_{x}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized, $g_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ encodes constraints and $\mathcal{S}_{x}=\{g_{x}(s)\leq 0\}$ is the set of feasible solutions, and $\Phi(x)$ is the instance’s optimal value. The paper assumes the availability of a primal-dual pair in the following form. A primal solution $\bar{s}\in\mathcal{S}_{x}$ provides a primal (upper) bound on the optimal value, i.e., $f_{x}(\bar{s})\geq\Phi(x)$. Primal solutions are obtained using either heuristics of exact algorithms. Conversely, a dual (lower) bound $\psi\in\mathbb{R}$ is guaranteed to be smaller than the optimal value, i.e., $\psi\leq\Phi(x)$. Dual bounds are obtained from relaxations and/or branch-and-bound algorithms, and provide a mathematical certificate on the quality of a primal solution $\bar{s}$. It is important to note that all modern global optimization algorithms return such a primal-dual pair. It is also possible to obtain a primal-dual pair using only machine learning (ML) models, by combining primal-feasible ML architectures Donti et al. (2021); Chen et al. (2023) with the dual lagrangian learning framework introduced in Tanneau and Hentenryck (2024).
General problem setting
Consider a random variable $X$ with support $\mathcal{X}\subseteq\mathbb{R}^{d}$, and an unknown Borel measurable function $f{:}\,\mathcal{X}\rightarrow\mathbb{R}$ that defines a target $Y\,{=}\,f(X)\,{\in}\,\mathbb{R}$. Let $\hat{B}^{l}$ and $\hat{B}^{u}$ be given lower and upper predictors for $f$, i.e.,
$\forall x\in\mathbb{R}^{d},\hat{B}^{l}(x)\leq f(x)\leq\hat{B}^{u}(x).$ (2)
These predictors inherently form a 100% prediction interval $\hat{B}^{l}(X),\hat{B}^{u}(X)]$ for $Y$. Recall that, in a constrained optimization context, upper and lower predictors are obtained from primal and dual information, respectively. The goal of this paper is to derive tighter prediction intervals by targeting a $(1-\alpha)$ marginal coverage for $Y$ (where $\alpha\,{\in}\,[0,1]$), rather than perfect coverage.
Formally, consider a dataset of $N$ independent and identically distributed (i.i.d.) observations $(X_{i}\,{,}\,Y_{i})_{i\in[N]}$, where $X_{i}$ is sampled from an underlying distribution $\mathcal{P}$, and $Y_{i}\,{=}\,f(X_{i}),\forall i\,{\in}[N]$. Following the standard procedure of the split CP framework Vovk et al. (2005), divide the $N$ samples into a training and a calibration set, denoted as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}},$ with corresponding indices $\mathcal{I}_{\text{train}},\mathcal{I}_{\text{cal}}$. Next, consider a new i.i.d. observation $X_{N+1}$, sampled from $\mathcal{P}$. For $\alpha\,{\in}\,[0,1]$, the goal is to refine the initial prediction interval $\tilde{C}(X_{N+1})=[\hat{B}^{l}(X_{N+1}),\hat{B}^{u}(X_{N+1})]$ into a (smaller) prediction interval $\hat{C}(X_{N+1})=[\hat{L}(X_{N+1}),\hat{U}(X_{N+1})]$ that satisfies the marginal coverage guarantee
$\mathbb{P}(Y_{N+1}\in\hat{C}(X_{N+1}))\geq 1-\alpha.$ (3)
Indeed, recall that, while $\tilde{C}(X_{N+1})$ has 100% coverage by construction, it may be too wide to derive meaningful insights. This motivates the need to derive smaller prediction intervals while ensuring reasonable marginal coverage. Then the goal can be restated as the optimization problem
$\min_{\hat{C}}\quad$ $\mathbb{E}_{X\sim\mathcal{P}}\left[\big|\hat{C}(X)\big|\right]$ (4a)
s.t. $\mathbb{P}(f(X)\in\hat{C}(X))\geq 1-\alpha,$ (4b)
where $\hat{C}$ maps $X\in\mathbb{R}^{d}$ to a subset of $\mathbb{R}$.
Remark 1. When multiple bounds $\{\hat{B}^{l}_{1},\ldots,\hat{B}^{l}_{B_{l}}\}$ and $\{\hat{B}^{u}_{1},\ldots,\hat{B}^{u}_{B_{u}}\}$ are provided, valid predictors $\hat{B}^{l}(x)$ and $\hat{B}^{u}(x)$ leveraging all this information effectively can be defined as
$\hat{B}^{l}(x)=\max\{\hat{B}^{l}_{1}(x),\ldots,\hat{B}^{l}_{B_{l}}(x)\},\quad\hat{B}^{u}(x)=\min\{\hat{B}^{u}_{1}(x),\ldots,\hat{B}^{u}_{B_{u}}(x)\}.$ (5)
2 Background
UQ is crucial for informed decision making, particularly in real-world applications where prediction sets are preferred over point estimates. CP, first proposed by Vovk et al. (2005), is a widely used distribution-free UQ method, valued for its finite-sample coverage guarantees and computational efficiency. This section presents standard CP techniques, including Split CP, Conformal Quantile Regression, and Nested CP. While these methods can be applied to the problem at hand (see Section 5 and Appendix LABEL:app:experiments:cp_baselines), unlike the proposed framework, they are not designed to explicitly leverage valid lower and upper bound predictors.
Split Conformal Prediction (SCP)
Vovk et al. (2005); Papadopoulos et al. (2002); Lei et al. (2018) is one of the most commonly used CP frameworks. Therein, a mean prediction model $\hat{f}$ is first trained using the training data $\mathcal{D}_{train}$. SCP then produces prediction intervals of the form
$\hat{C}_{\text{SCP}}(x)=\left\{y\,\middle|\,\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}}\leq\hat{s}(x,y)\leq\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}\right\},$ (6)
where $\hat{s}\,{:}\,\mathbb{R}^{d}{\times}\mathbb{R}\,{\rightarrow}\,\mathbb{R}$ is a so-called conformity score, and $\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}},\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}$, denote the empirical $\frac{\alpha}{2}$ and $1{-}\frac{\alpha}{2}$ quantiles of the scores $\{\hat{s}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$. Under the exchangeability assumption, the resulting prediction intervals have valid marginal coverage Vovk et al. (2005), i.e.,
$\mathbb{P}_{X_{N+1}\sim\mathcal{P}}\left(Y_{N+1}\in\tilde{C}_{\text{SCP}}(X_{N+1})\right)\geq 1-\alpha.$ (7)
Common choices of conformity scores include the residual score $\hat{s}(x,y)\,{=}\,y{-}\hat{f}(x)$ and the absolute residual score $\hat{s}(x,y)\,{=}\,|y{-}\hat{f}(x)|$; the reader is referred to Angelopoulos and Bates (2021); Oliveira et al. (2024) for a more exhaustive review of SCP and conformity scores. The SCP methodology can be applied to the paper’s setting by replacing $\hat{f}$ in the above construction with either the lower ($\hat{B}^{l}$) or upper bound ($\hat{B}^{u}$) predictor. However, this fails to jointly utilize information from both bounds.
Conformal Quantile Regression (CQR)
To alleviate SCP’s lack of local adaptivity, CQR Romano et al. (2019) combines quantile regression models with a conformalization procedure: First, CQR trains quantile regression models to predict the quantiles $\alpha_{\text{lo}}$ and $\alpha_{\text{hi}}$. Then, conformalize as follows
$\hat{C}^{\text{CQR}}(x)=\left[\hat{q}_{\alpha_{\text{lo}}}(x){-}\hat{Q}^{\text{CQR}}_{1-\alpha},\ \hat{q}_{\alpha_{\text{hi}}}(x){+}\hat{Q}^{\text{CQR}}_{1-\alpha}\right],$ (8)
where $\hat{q}_{\alpha_{\text{lo}}}(x),\hat{q}_{\alpha_{\text{hi}}}(x)$ are the predicted lower and upper quantiles given input data $x$, and $\hat{Q}^{\text{CQR}}_{1-\alpha}$ is the $1{-}\alpha$ quantile of the conformity scores $\{\hat{s}^{\text{CQR}}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$, defined as
$\hat{s}^{\text{CQR}}(x,y)=\max\left(\hat{q}_{\alpha_{\text{lo}}}(x){-}y,\ y{-}\hat{q}_{\alpha_{\text{hi}}}(x)\right).$ (9)
The CQR method could be adapted to the setting at hand by approximating $\hat{q}_{\alpha_{\text{lo}}},\hat{q}_{\alpha_{\text{hi}}}$ with $\hat{B}^{l},\hat{B}^{u}$ in the above derivation. Unlike the standard CQR, which requires fitting auxiliary quantile regression models whose accuracy dictates performance (and may require large training datasets in high-dimensional settings), $\hat{B}^{l},\hat{B}^{u}$ from the constrained optimization formulation offer key advantages. These bounds are often directly computable from the base model (e.g., once a feasible primal model is constructed, an upper bound can be computed by evaluating the objective at that solution), making them available at minimal cost. They also inherently guarantee perfect coverage, regardless of training data size. This paper focuses on effectively leveraging these readily available and valid bounds.
Nested Conformal Prediction (NCP)
Gupta et al. (2022) introduces a unifying CP framework by nested prediction sets. NCP considers a family of nested prediction intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}\subseteq\mathbb{R}}$, i.e.,
$\forall x\in\mathcal{X},\forall t\leq t^{\prime},\hat{C}_{t}(x)\subseteq\hat{C}_{t^{\prime}}(x),$ (10)
and such that $\hat{C}_{\inf(\mathcal{T})}=\emptyset$ and $\hat{C}_{\sup(\mathcal{T})}=\mathbb{R}$. Then, prediction intervals are of the form $\hat{C}^{\text{NCP}}(x)=\hat{C}_{\tau}(x)$, where
$\tau=\inf_{t\in\mathcal{T}}\left\{t\,\middle|\,\sum_{i\in\mathcal{I}_{\text{cal}}}\mathbf{1}_{\hat{C}_{t}(X_{i})}(Y_{i})\geq(1{-}\alpha)(1{+}|\mathcal{I}_{cal}|)\right\}$ (11)
is computed from the calibration set. This framework encompasses all types of conformity scores, while maintaining the theoretical guarantees of standard CP Gupta et al. (2022). For instance, the CQR construction can be cast in the NCP framework by considering
$\hat{C}_{t}(x)=\left[\hat{q}_{\frac{\alpha}{2}}(x)-t,\ \hat{q}_{1-\frac{\alpha}{2}}(x)+t\right].$ (12)
To ensure clarity, the remainder of the paper is presented using the NCP framework. Namely, each CP methodology is presented by stating the corresponding family of nested prediction sets; prediction intervals are then constructed using the NCP approach and Eq. (11).
Other Related Methods
CP techniques have been use to analyze the precision of sketching algorithms, which are increasingly vital tools for handling massive datasets in modern machine learning Broder and Mitzenmacher (2004); Goyal et al. (2012); Cormode et al. (2018); Zhang et al. (2014). Sesia et al. (2023) construct prediction intervals for the frequency of a queried object based on a sketch Charikar et al. (2002). Therein, a deterministic upper bound $\hat{B}^{u}$ provided by Cormode and Muthukrishnan (2005) is combined with a trivial lower bound $\hat{B}^{l}=0$ (reflecting the non-negativity of counts) to construct nested intervals. The construction, referred to as “SFD CP” in this paper, is defined as:
$\hat{C}^{\text{SFD}}_{t}(x)=\big[\max\{0,\hat{B}^{u}(x){-}t\},\min\{\hat{B}^{u}(x),t\}\big].$ (13)
Sesia et al. (2023) also introduces an adaptive version of this construction, designed to better account for heteroscedasticity in residuals. The CPUL framework generalizes this approach by combining information from $\hat{B}^{l},\hat{B}^{u}$ in multiple ways; see Section 3.2.
Another related line of work leverages information from multiple models fitted on the training set. For example, Liang et al. (2024); Yang and Kuchibhotla (2024) focus on selecting the model that produces the most efficient prediction intervals. Extending these ideas, the CPUL framework reinterprets the problem in the context of multiple models, simultaneously utilizing both upper and lower bounds to maximize the utility of the available information.
3 CP With Upper And Lower Bounds
(a) Illustration of the construction of prediction intervals in CPUL: (a) $\hat{C}^{\text{ll}}_{0}$, (b) $\hat{C}^{\text{uu}}_{0}$, (c) $\hat{C}^{\text{lu}}_{0}$, (d) $\hat{C}^{\text{ul}}_{0}$. Each family of prediction intervals is conformalized following the NCP framework (see Eq. (11)).
(b) Illustration of Paradoxical Miscoverage (motivation for OMLT): using a constant offset ($\pm t$) in the NCP construction results in an empty prediction interval where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is small.
Figure 1: Illustration of CPUL-OMLT construction
This section presents CPUL, designed to exploit the strong properties of $\hat{B}^{l},\hat{B}^{u}$. A key limitation of traditional CP methods like SCP and CQR is not leveraging that $\hat{B}^{l},\hat{B}^{u}$ yield valid lower and upper bounds on the target variable. Additionally, these methods do not necessarily integrate information from both models nor are they inherently designed to account for heteroskedastic residuals between these bounds (e.g., $\hat{B}^{l}$ might provide accurate estimates while $\hat{B}^{u}$ performs poorly). The results in Section 5 further validate these observations.
3.1 Exploiting valid bounds in NCP
Recall that $\hat{B}^{l},\hat{B}^{u}$ provide valid lower and upper bounds on the target variable, i.e., $\hat{B}^{l}(X)\,{\leq}\,Y\,{\leq}\,\hat{B}^{u}(X)$ always holds. This suggests a simple procedure for strengthening $\hat{C}$ without loss of coverage (as shown by Proposition 1) into
$\tilde{C}(x)=\hat{C}(x)\cap[\hat{B}^{l}(x),\hat{B}^{u}(x)].$ (14)
Proposition 1.
[…]
$\hat{C}(x)\,{\cap}\,\big[\hat{B}^{l}(x),\hat{B}^{u}(x)\big],\forall x\in\mathcal{X}$. Then, $\mathbb{P}\big(Y_{N+1}\,{\in}\,\tilde{C}(X_{N+1})\big)=\mathbb{P}\big(Y_{N+1}\,{\in}\,\hat{C}(X_{N+1})\big)=1{-}\alpha.$
It is important to note that Proposition 1 holds irrespective of how the original prediction interval $\hat{C}$ is obtained. Theorem 1 shows that, if the calibration step is performed using the NCP framework, then there is no loss of performance whether the strengthening is performed before or after calibration.
Theorem 1.
Consider a family of nested prediction sets $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, where $\mathcal{T}\,{\subseteq}\,\mathbb{R}$, and let $\hat{\tau}$ be obtained following the NCP calibration step as per Eq. (11). Next, define the family of nested strengthened intervals $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$, where $\forall x\in\mathbb{R}^{d},\forall t\in\mathcal{T},\tilde{C}_{t}(x)=\hat{C}_{t}(x)\cap\big[\hat{B}^{l}(x),\hat{B}^{u}(x)\big],$ and let $\tilde{\tau}$ be obtained from Eq. (11). Then, $\hat{\tau}=\tilde{\tau}$ and $\forall x\in\mathbb{R}^{d},$
$\hat{C}_{\hat{\tau}}(x)\cap[\hat{B}^{l}(x),\hat{B}^{u}(x)]$ $=\tilde{C}_{\tilde{\tau}}(x).$ (15)
Theorem 1 allows to decouple the calibration step from the strengthening. Thereby, one can perform the calibration step without access to $\hat{B}^{l}$ or $\hat{B}^{u}$, without any loss of coverage. This is particularly valuable when $\hat{B}^{l}$ or $\hat{B}^{u}$ are not available during training/calibration due to, e.g., privacy concerns.
3.2 The CPUL approach
The proposed CPUL framework combines several steps to fully exploit the knowledge that $\hat{B}^{l}$ and $\hat{B}^{u}$ provide valid bounds on the target variable, and to account for the heteroskedasticity of their residuals. The algorithm is presented in the NCP framework and, for ease of reading, the presentation omits the strengthening of prediction intervals described in Eq. (14), i.e., strengthening is always performed implicitly. Recall that, by Theorem 1, this does not affect the conformalization procedure.
First define the residuals $\hat{r}^{l}\,{=}\,Y\,{-}\,\hat{B}^{l}(X)$ and $\hat{r}^{u}\,{=}\,Y\,{-}\,\hat{B}^{u}(X)$, and denote by $\hat{Q}^{l}_{\beta}$, $\hat{Q}^{u}_{\beta}$ the $\beta$ quantiles of $\hat{r}^{l},\hat{r}^{u}$, evaluated on the training set. Then define
$L^{l}(x)$ $=\hat{B}^{l}(x)+\hat{Q}^{l}_{\frac{\alpha}{2}},$ (16a)
$L^{u}(x)$ $=\hat{B}^{u}(x)+\hat{Q}^{u}_{\frac{\alpha}{2}},$ (16b)
$U^{l}(x)$ $=\hat{B}^{l}(x)+\hat{Q}^{l}_{1-\frac{\alpha}{2}},$ (16c)
$U^{u}(x)$ $=\hat{B}^{u}(x)+\hat{Q}^{u}_{1-\frac{\alpha}{2}}.$ (16d)
Note that $L^{l}$, $L^{u}$, $U^{l}$, $U^{u}$ do not provide valid lower nor upper bounds on the target variable ($Y$). The motivation for adjusting the initial predictors $\hat{B}^{l},\hat{B}^{u}$ using quantiles $\hat{Q}^{l},\hat{Q}^{u}$ is to account for the heteroskedascitity of $\hat{r}^{l},\hat{r}^{u}$. Also note that the construction in Eq. (16) only requires evaluating the residuals $\hat{r}^{l},\hat{r}^{u}$ on the training set, and extracting their quantiles.
Next, $L^{l}$, $L^{u}$, $U^{l}$, $U^{u}$ are combined to form four families of nested prediction intervals $\{\hat{C}^{\text{ll}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{lu}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{ul}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{uu}}_{t}\}_{t\in\mathbb{R}}$ as follows:
$\hat{C}^{\text{ll}}_{t}(x)$ $=\big[L^{l}(x)-t,U^{l}(x)+t\big],$ (17a)
$\hat{C}^{\text{lu}}_{t}(x)$ $=\big[L^{l}(x)-t,U^{u}(x)+t\big],$ (17b)
$\hat{C}^{\text{ul}}_{t}(x)$ $=\big[L^{u}(x)-t,U^{l}(x)+t\big],$ (17c)
$\hat{C}^{\text{uu}}_{t}(x)$ $=\big[L^{u}(x)-t,U^{u}(x)+t\big].$ (17d)
This construction is illustrated in Figure 1(a). Each family is then conformalized using the NCP procedure on the calibration set, and the model with smallest width is selected.
Algorithm 1 summarizes the proposed CPUL method, which proceeds as follows. First, the dataset $\mathcal{D}$ is split into training and calibration sets (line 1). Then, lower and upper bound predictors (i.e., $\hat{B}^{l},\hat{B}^{u}$) are trained using the training set (line 2), after which quantiles of the corresponding residuals are evaluated on the training set (line 3). Next, four families of nested intervals are constructed following Eq. (17), each of which is calibrated using the NCP framework (Eq. (11)) and the calibration set $\mathcal{D}_{\text{cal}}$ (line 4). The final step of the algorithm (line 5) selects, among the four models $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, the one with smallest average width over the calibration set (as an estimator of $\mathbb{E}\big|\hat{C}(X_{N+1})\big|$). It is important to note that the model selection step is not performed on a per-sample basis. Rather, Algorithm 1 selects one variant among $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, which is then used across the entire test set. For instance, if $\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ is selected, then $\hat{C}^{**}\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ and the CPUL prediction interval is $\hat{C}^{**}(X_{N+1})\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}(X_{N+1})$. Theorem 2 provides theoretical guarantees on the coverage of $\hat{C}^{**}$.
Theorem 2.
Assume that $\{(X_{1},Y_{1}),...,(X_{N+1},Y_{N+1})\}$ are i.i.d. samples, and let $\hat{C}^{**}$ be the CPUL model selected by Algorithm 1. Define $N_{\text{cal}}=|\mathcal{D}_{\text{cal}}|$ and $\eta=\sqrt{\nicefrac{{\log(8)}}{{2}}}+\nicefrac{{1}}{{3}}$. Then
$\mathbb{P}\left(Y_{N+1}\in\hat{C}^{**}(X_{N+1})\right)\geq\frac{1{+}N_{\text{cal}}}{N_{\text{cal}}}(1{-}\alpha)-\frac{\eta}{\sqrt{N_{\text{cal}}}}.$ (18)
Proof.
The result follows directly from Theorem 1 from Yang and Kuchibhotla (2024). ∎
3.3 Relation to other methods
Several parallels can be drawn between CPUL and existing CP methodologies. For instance, the use of residual quantiles $\hat{Q}^{l},\hat{Q}^{u}$, when constructing the nested prediction sets $\hat{C}^{\text{ll}},\hat{C}^{\text{uu}}$, is related to the Split CP method. Indeed, applying SCP to $\hat{B}^{l}$ or $\hat{B}^{u}$ yields conformal prediction intervals (see Eq. (6)) that closely resemble the structure of $\hat{C}^{\text{ll}}$ and $\hat{C}^{\text{uu}}$. Therefore, given that CPUL also exploits the $\hat{C}^{\text{lu}}$ and $\hat{C}^{\text{ul}}$ variants, one should expect CPUL to consistently outperform either SCP baseline.
The $\hat{C}^{\text{ul}}$ construction is also related to the adaptive approach followed in Sesia et al. (2023). The main difference between the two is the use of empirical quantiles $\hat{Q}^{l},\hat{Q}^{u}$ in CPUL, which can be obtained much more efficiently, compared to training additional quantile regression models as in Sesia et al. (2023). The latter approach provides finer local adaptivity, albeit at a higher computational cost.
Algorithm 1 CPUL
Input: $\mathcal{D}\,{=}\,({X_{i}},{Y_{i}})_{i=1}^{N}$, $\alpha\,{\in}\,[0,1]$.Output: Selected model $\hat{C}^{**}$.
1:  Split $\mathcal{D}$ into training and calibration sets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}}$
2:  Train lower/upper bound models $\hat{B}^{l},\hat{B}^{u}$ using $\mathcal{D}_{\text{train}}$
3:  Compute empirical quantiles $\hat{Q}^{l}_{\frac{\alpha}{2}}$, $\hat{Q}^{l}_{1-\frac{\alpha}{2}}$, $\hat{Q}^{u}_{\frac{\alpha}{2}}$, $\hat{Q}^{u}_{1-\frac{\alpha}{2}}$ on the training set $\mathcal{D}_{\text{train}}$
4:  Form $\{\hat{C}^{\text{ll}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{lu}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{ul}}_{t}\}_{t\in\mathbb{R}}$, $\{\hat{C}^{\text{uu}}_{t}\}_{t\in\mathbb{R}}$ as per Eq. (17), and compute $\tau^{\text{ll}}$, $\tau^{\text{lu}}$, $\tau^{\text{ul}}$, $\tau^{\text{uu}}$ using Eq. (11).
5:  Select model with smallest width on calibration set
[…]
4 Optimal Minimal Length Threshold
A key limitation of CPUL, like most well-known CP methods, is its tendency to produce overly narrow prediction intervals, particularly when the initial bounds $\hat{B}^{u}(x)$ and $\hat{B}^{l}(x)$ are tight. The length of the prediction intervals in Eq. (16) scales with the magnitude of $\hat{B}^{u}(x)-\hat{B}^{l}(x)$. Consequently, when the bounds are extremely tight, the constant adjustment of $t$ in the NCP framework becomes disproportionately large, leading empty prediction intervals and, in turn, under-coverage in such regions. Figure 1(b) illustrates this issue in the CPUL-lu setting. After calibrating the initial bounds the area around $x_{0}$ remains uncovered due to the empty intersection of the calibrated intervals. This behavior results in inefficient prediction intervals, as the method should over-cover the less confident regions (i.e., areas where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is large). Hence, the regions with tightest initial bounds become most vulnerable and become under-covered. Sesia and Candès (2020) proposed a variant, CQR-r, with $\hat{C}_{t}^{\text{CQR-r}}(x)=\left[\hat{L}_{t}^{\text{CQR-r}}(x),\hat{U}_{t}^{\text{CQR-r}}(x)\right],\ \text{for }x\in\mathcal{X},$ where
$\hat{L}_{t}^{\text{CQR-r}}(x)=$ $\hat{q}_{\alpha/2}(x)-t\left(\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)\right);$
$\hat{U}_{t}^{\text{CQR-r}}(x)=$ $\hat{q}_{1-\alpha/2}(x)+t\left(\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)\right).$
The design of Sesia and Candès (2020) scales $t$ by $\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)$, which mitigates the disproportionate reduction in prediction interval size. However, as noted in Sesia and Candès (2020), its overall efficiency is worse than the original fixed-length version (12). Thus, it remains unclear whether such scaling, despite being adaptive to the initial tightness of the base models, is preferred over the fixed-length adjustment $t$. Section 5 includes a CQR-r adaptation tailored to the setting of this paper for comparison.
To address this paradoxical miscoverage, the paper proposes the optimal minimal length threshold method (OMLT) as an alternative to the relative scaling approach used in Sesia and Candès (2020). The core idea of OMLT consists in introducing a threshold $\ell\geq 0$, representing the minimum allowed length for a prediction interval. OMLT identifies regions where the given $\hat{B}^{u}$ and $\hat{B}^{l}$ are tightest, where even minor shrinking during calibration poses a high risk of significant undercoverage below the desired level of $1-\alpha$. The design of OMLT retains the standard fixed-length adjustment of $t$ and introduces a threshold for the minimal allowed prediction interval length, effectively marking the boundary of high-risk regions.
Consider a family of nested intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$ satisfying the NCP assumptions, and define
$\kappa_{\ell}(x)=\inf_{t\in\mathcal{T}}\big\{t\ \big|\ \ell\leq|\hat{C}_{t}(x)|\big\},$ (19)
from which a new family $\{\bar{C}_{t}\}_{t\in\mathcal{T}}$ is constructed as
$\bar{C}_{\ell,t}(x)=\left\{\begin{array}[]{ll}\hat{C}_{t}(x)&\text{if }(\Delta(x)\,{\geq}\,\ell)\wedge(t\,{>}\,\kappa_{\ell}(x))\\ \hat{C}_{\kappa_{\ell}(x)}(x)&\text{if }(\Delta(x)\,{\geq}\,\ell)\wedge(t\,{\leq}\,\kappa_{\ell}(x))\\ {[\hat{B}^{l}(x),\hat{B}^{u}(x)]}&\text{if }(\Delta(x)\,{\leq}\,\ell)\end{array}\right$. (23)
where $\Delta(x)\,{=}\,\hat{B}^{u}(x)\,{-}\,\hat{B}^{l}(x)$. It is easy to verify that $\{\bar{C}_{\ell,t}\}_{t\in\mathcal{T}}$ is a nested family, therefore satisfying the coverage guarantee given in (11). The minimum length threshold $\ell$ may be specified by the user as a hyper-parameter, obtained as the solution of the optimization problem
$\min_{\ell\geq 0,t\in\mathcal{T}}\quad$ $\mathbb{E}_{X}\left[\left|\bar{C}_{\ell,t}(X)\right|\right]$ (24a)
s.t. $\mathbb{P}\left(f(X)\in\bar{C}_{\ell,t}(X)\right)\geq 1-\alpha.$ (24b)
The key idea underlying OMLT is that the size of the prediction interval should not be smaller than $\ell$, which prevents under-coverage when a prediction interval is too small. The only exception is for the tightest regions, i.e., when $\hat{B}^{u}(x)-\hat{B}^{l}(x)\leq\ell$, in which case there is no need to enlarge $\bar{C}_{\ell,t}(x)$ beyond $\ell$. The original upper and lower bounds can be confidently relied upon, as $\hat{B}^{l}(x),\hat{B}^{u}(x)]$ already provides high-quality coverage.
Since (24) reduces to the original problem in (4) when $\ell=0$ for arbitrary $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, the formulation of CPUL-OMLT is designed to produce prediction intervals whose expected length is generally not greater than that of CPUL, at the same $1-\alpha$ coverage. Moreover, OMLT has the potential to extend beyond this setting and reduce the average prediction interval length in cases where disproportionate prediction calibration is observed. For details of implementation, refer to Section LABEL:sec:CPUL_and_OMLT_implementation_details.
5 Experiments
The performance of CPUL and CPUL-OMLT are evaluated to demonstrate their ability to achieve valid coverage while producing narrower prediction intervals. The experiments focus on UQ for the optimal value of economic dispatch problems, detailed in Appendix LABEL:sec:background:optimization. UQ is conducted given lower and upper bounds derived from primal-dual optimization proxy models Chen et al. (2023); Qiu et al. (2024); Klamkin et al. (2024); Chen et al. (2024). Namely, the dual proxy provides valid lower bounds ($\hat{B}^{l}$), and the primal proxy provides valid upper bounds $\hat{B}^{u}$). Further details on these optimization proxies are provided in Appendices LABEL:sec:background:optimization and LABEL:sec:background:proxies.
Table 1: Performance Comparison of CP Methods ($\alpha=10\%$)
89_pegase 118_ieee 1354_pegase
UQ Method PICP (%) Size (%) PICP (%) Size (%) PICP (%) Size (%)
$\hat{B}^{l},\hat{B^{u}}$] 100.0 (0.00) 0.410 (0.016) 100.0 (0.00) 0.281 (0.085) 100.0 (0.00) 3.949 (3.807)
Split CP w/ $\hat{B}^{l}$ 90.15 (0.60) 0.197 (0.007) 90.34 (0.63) 0.206 (0.146) 89.60 (0.54) 1.308 (1.160)
Split CP w/ $\hat{B}^{u}$ 90.02 (0.54) 0.205 (0.007) 90.00 (0.50) 0.105 (0.004) 89.40 (0.49) 1.570 (0.853)
SFD CP 91.23 (0.56) 0.188 (0.007) 90.10 (0.50) 0.135 (0.013) 90.22 (0.27) 1.456 (0.609)
CQR 91.40 (1.80) 0.389 (0.018) 90.08 (2.75) 0.190 (0.147) 90.13 (0.43) 3.501 (3.939)
CQR-r 90.71 (0.72) 0.332 (0.020) 90.88 (1.33) 0.180 (0.167) 91.57 (1.32) 3.271 (3.997)
CPUL (ours) 91.23 (0.56) 0.188 (0.007) 90.02 (0.46) 0.105 (0.004) 89.65 (0.56) 1.306 (1.162)
CPUL-OMLT (ours) 90.28 (0.51) 0.187 (0.008) 90.01 (0.48) 0.103 (0.007) 89.69 (0.49) 1.037 (0.890)
    For each dataset, the three shortest intervals are colored blue, while the three largest intervals are colored red.
Datasets
Experiments are performed over three datasets: 118_ieee University of Washington, Dept. of Electrical Engineering (1999), 1354_pegase Fliscounakis et al. (2013), and 89_pegase Fliscounakis et al. (2013), corresponding to power grids of different sizes. For each dataset, samples are randomly shuffled and split 10 times. Dataset details are provided in Section LABEL:data123. Evaluation metrics reported are the mean and standard deviation across these 10 runs.
Baselines
The performance of CPUL-OMLT is compared against several CP baselines. Recall that, unless specified otherwise, all prediction intervals are strengthened as per Eq. (14). The CP baselines include: Split CP (Eq. 6), using either $\hat{B}^{l}$ or $\hat{B}^{u}$ as base predictors; CQR and its variant CQR-r, wherein $\hat{B}^{l}$ and $\hat{B}^{u}$ are treated as the initial lower and upper quantile regressors in the CQR construction; an adapted SFP CP approach of Sesia et al. (2023). Note that the latter matches the $\hat{C}^{\text{ul}}$ construction. Additional implementation details are provided in Appendix LABEL:app:experiments:cp_baselines.
Evaluation Metrics
All methods are evaluated on a held out test set $\mathcal{D}_{\text{test}}\,{=}\,(X_{i},Y_{i})_{i\in\mathcal{I}_{\text{test}}}$, using two evaluation criteria: coverage and interval length. The former is evaluated via the Prediction Interval Coverage Percentage (PICP), which measures the proportion of true values contained within the predicted intervals on the test set, i.e., $\text{PICP}=\frac{1}{|\mathcal{I}_{\text{test}}|}\sum_{i\in\mathcal{I}_{\text{test}}}\mathbf{1}_{\hat{C}(X_{i})}(Y_{i}).$
At a given confidence level $1\,{-}\,\alpha$, shorter interval lengths are considered more desirable. This is captured through expected normalized length of prediction interval $\hat{C}$, defined as $\hat{E}_{X}(\hat{C})=|\mathcal{I}_{\text{test}}|^{-1}\sum_{i\in\mathcal{I}_{\text{test}}}\big(\left|Y_{i}\right|^{-1}|\hat{C}(X_{i})|\big)$, where $Y_{i}$ and $\hat{C}(X_{i})$ denote the (true) optimal value and the prediction interval for sample $i$. The paper considers the scaled interval length $|\hat{C}(X_{i})|/|Y_{i}|$, expressed as a percentage, rather than absolute interval length $|\hat{C}(X_{i})|$, to account for the possibly large range of values taken by $Y$. This metric is routinely used in the optimization literature Chen et al. (2023); Vivas et al. (2020); Wang et al. (2009), where it is referred to as optimality gap. All metrics are reported as the mean and standard deviation over 10 runs.
Experiment Results
Figure 2: Performance comparison across datasets (Note: in 89_pegase, SFD CP overlaps with CPUL and in 118_ieee, Split CP w/ $\hat{B}^{l}$ overlaps with CPUL.)
Table 1 and Figure 2 present the numerical performance of the various methods. Table 1 presents, for every dataset, the average and standard deviation of each method’s interval length and coverage, both expressed as a percentage. The table reports results for $\alpha=10\%$; additional $\alpha$ values are reported in Table LABEL:table_full:results in Appendix LABEL:sec:Additional_experiment_results. Figure 2 displays the interaction between each method’s interval length and coverage, across a broader range of $\alpha$ values. This provides a more global view of each method’s overall performance.
The results in Table 1 demonstrate that all CP methods achieve satisfactory coverage levels, i.e., test coverage is typically close to $1-\alpha$, which is expected given that all methods are properly calibrated. Moreover, all prediction intervals yield a substantial reduction in size compared to the original prediction $\hat{B}^{l},\hat{B}^{u}]$. For instance, a reduction of only 10% in coverage (i.e., 90% prediction intervals) can yield up to a two- and three-fold reduction in interval size. This demonstrates the value of using UQ techniques to provide more actionable information to practitioners.
Furthermore, CPUL and CPUL-OMLT consistently achieve state-of-the-art performance, across datasets and target coverage. On the other hand, CQR and CQR-r exhibit the worst performance overall, likely due to their inability to properly address heteroskedasticity. This is most evident in Figure 2. On the 1354_pegase dataset (which corresponds to a larger power grid) CPUL-OMLT achieves the smallest interval size. In particular, CPUL-OMLT produces prediction intervals whose witdth is about 10% smaller than the second-best method, CPUL. These results highlight the important of the model selection procedure in CPUL.
Figure 2 demonstrates that, while methods like Split CP and SFD-CP deliver good results occasionally, their performance is highly variable across different datasets and confidence levels. For instance, although SFD-CP is among the best-performing methods on 89_pegase, its performance on 1354_pegase is significantly worse than CPUL when coverage is close to 100% (corresponding to small values of $\alpha$). Similarly, Split CP using $\hat{B}^{u}$ is among the worst performers on the 118_IEEE dataset, but among the top performers on the 1354_pegase dataset. Such variability in performance further reinforce the value of model selection, which is most evident on the 1354_pegase dataset, Thereby, methods such as SCP with $\hat{B}^{l}$ and SFP CP perform well when the coverage is between 85% and 90%, but are out-performed by SCP with $\hat{B}^{u}$ when coverage is above 90%. In contrast, CPUL and CPUL-OMLT perform well across the entire coverage range. This is explained by the fact that CPUL and CPUL-OMLT select the best construction for each target coverage $\alpha$.
Finally, while CPUL and CPUL-OMLT perform similarly on the 89_pegase and 118_IEEE datasets, CPUL-OMLT offers a clear improvement on the 1354_pegase dataset, especially for coverage levels above 90% (see Figure 2). A more granular analysis of each model’s behavior reveals that CPUL’s performace is worse on samples where $\hat{B}^{l},\hat{B}^{u}]$ is small. Namely, CPUL achieves a coverage of only about 50% across the 5% of test samples with smallest initial interval $\hat{B}^{l},\hat{B}^{u}]$. Recall that these samples correspond to regions where the given bounds $\hat{B}^{l},\hat{B}^{u}$ are the most accurate, which demonstrates the issue of paradoxical miscoverage (see Section 4 and Figure 1(b)). CPUL-OMLT effectively mitigates this issue, which results in better coverage and smaller interval length overall.
6 Conclusion
This paper introduced CPUL-OMLT, a novel CP mechanism designed for the settings that deterministic lower and upper bounds on the target variable are available. By integrating multiple interval construction strategies within the NCP framework, CPUL effectively leverages the structure of these bounds to improve efficiency. The OMLT mechanism leverages the strong condition that tight initial bounds already provide highly precise confidence intervals. Failure to account for this often leads to undercoverage in such regions, a challenge that existing methods struggle to address. Experimental validations were conducted on optimization problems across three power systems datasets, demonstrating that the proposed approach consistently outperforms the traditional CP methods by providing better efficiency. The proposed CPUL-OMLT method provides state-of-the-art efficiency and remains consistent across different datasets, highlighting its robustness and practical relevance.
Future work could extend CPUL-OMLT beyond the efficient marginal coverage studied here. Potential directions include enforcing stronger conditional coverage to incorporate the already computed bounds $\hat{B}^{l},\hat{B}^{u}$, integrating with optimization algorithms like branch and bound to directly boost downstream task performance, and adapting the method for high-dimensional/structured data to increase applicability in complex settings like energy systems.